the length of a rectangle is 4 times the width. The perimeter is 90 cm. how can i set up this equation
Answer by checkley77(12844)

L=4W
2L+2W=90
2*4W+2W=90
8W+2W=90
10W=90
W=90/10
W=9 ANS. FOR THE WIDTH.
L=4*9=36 ANS. FOR THE LENGTH.
PROOF:
2*36+2*9=90
72+18=90
90=90